# Attention-Aided MMSE for OFDM Channel Estimation: Learning Linear Filters with Attention

### v2

@@ I Introduction
Orthogonal frequency division multiplexing (OFDM) is a foundational waveform in modern wireless communication systems. OFDM was initially adopted in earlier standards such as Wi-Fi (IEEE 802.11), WiMAX, and was later introduced into the 3GPP-based cellular ecosystem with LTE. This adoption paved the way for OFDM’s continued prominence in 5G [1, 2] and its expected role in future 6G technologies [3].
@@
One of the critical problems in OFDM is channel estimation, which is essential for enabling coherent detection. The goal of OFDM channel estimation is to accurately estimate the channel state information for each subcarrier using noisy pilot symbols. A classical solution for the OFDM channel estimation involves signal processing (SP)-based approaches, including least square (LS) and minimum mean-squared error (MMSE) estimation methods [4, 5]. LS is simple to implement, but its performance is limited because it cannot incorporate prior OFDM channel statistics. MMSE, in contrast, requires second-order statistics, such as noise variance and correlation across subcarriers, which are difficult to obtain accurately in practical scenarios [6].
@@
Beyond traditional SP-based approaches, sensing-aided channel estimation has emerged within the integrated sensing and communication (ISAC) frameworks [7, 8]. These methods leverage communication signal echoes to extract sensing parameters, which then serve as prior information for channel refinement. Despite showing promise in exploiting communication-sensing synergy, these approaches require additional computational resources for sensing processing.
@@
Recently, inspired by the remarkable success of deep neural networks (DNNs) across various engineering fields, DNN-aided channel estimation for OFDM has gained considerable attention. DNNs, as pre-trained neural network models, can effectively capture complex data relationships without explicit prior statistical knowledge. Leveraging this data-driven approach, several prior works have explored DNN-aided OFDM channel estimation strategies. In [9], an OFDM channel slot was treated as image data. Based on this, a super-resolution network and a denoising neural network were simultaneously employed to estimate the OFDM channel (i.e., ChannelNet). In [10], using a similar modeling approach to [9], ReEsNet [10] was proposed, where a residual learning-based DNN was used to reconstruct the OFDM channel. This method achieves lower computational cost and higher accuracy than [9]. In [11], a residual convolutional network combined with bilinear interpolation was proposed for OFDM channel estimation, offering enhanced flexibility to varying pilot patterns and reduced complexity. Beyond deterministic models, [12] optimized a deep generative model using compressive pilot measurements for high-dimensional channel estimation. This method leverages learned priors to reduce pilot overhead and surpass traditional sparse recovery techniques.
@@
In particular, with the advent of the Attention Transformer architecture [13], some recent works investigated Transformer models as a compelling solution for wireless channel estimation problems, owing to their representational capacity. In [14], super-resolution techniques were integrated with an Attention-based encoder-decoder framework to enhance OFDM channel estimation accuracy. In [15], Channelformer was developed, where an Attention-based encoder-decoder neural architecture was designed to incorporate input precoding. A common structure in prior studies using DNNs for channel estimation [9, 10, 11, 12, 14, 15] is the use of a pre-trained neural network that directly maps noisy pilot signals to the estimated channel matrix, as illustrated in the left panel of Fig. 1.
@@
  • Computational overhead: Typically, DNN-based models involve a large number of parameters (approximately $53,000$ for ReEsNet [10] and $110,000$ for Channelformer [15]), leading to significant computational and memory overhead per channel estimation. Furthermore, several nonlinear operations are used (e.g., ReLU [16], tanh [17], ELU [18], and pooling [19]), limiting hardware acceleration effectiveness. These nonlinearities, although essential for model expressiveness, may introduce additional computational overhead and pose challenges related to inference complexity and practical implementation.
  • Limited interpretability: Existing DNN-based approaches are typically applied directly to OFDM channel estimation in a purely data-driven manner, without incorporating domain knowledge derived from mathematical models. However, studies such as [20, 21] showed that integrating domain-specific insights into neural network design can lead to notable performance gains over purely data-driven methods. This implies that considerable potential remains to further push performance boundaries by leveraging domain-specific insights.
@@
To address the challenges of purely data-driven approaches, recent research has actively explored the concept of model-based DNNs [22]. The core idea is to leverage a hybrid method that combines classical SP-based approaches and data-driven DNN-based approaches, thereby harnessing the advantages of both methodologies, i.e., powerful learning capability and mathematical interpretability. For instance, [23] proposed KalmanNet, where the Kalman gain matrix is learned using recurrent neural network architectures such as long short-term memory (LSTM) [24] or gated recurrent units (GRU) [25], while retaining the recursive update structure of the classical Kalman filtering algorithm [26]. Thanks to its hybrid design, KalmanNet [23] effectively address the state estimation problem, even with partially known state models. Subsequently, several extensions were proposed, including [27, 28, 29, 30]. In parallel, model-based DNN approaches have been widely applied to various tasks, such as multiple-input multiple-output (MIMO) beamforming design [31, 32], and MIMO detection [21, 33]. However, to the best of our knowledge, no model-based DNN framework has been explored for OFDM channel estimation. This work addresses this notable gap by devising a novel channel estimation method that leverages not only the learning capability of DNNs but also domain knowledge from classical SP-based approaches.
@@
Via numerical simulations across diverse 3GPP TDL scenarios, we show that the proposed A-MMSE and RA-A-MMSE significantly outperform traditional SP-based approaches (e.g., LS and MMSE) and recent DNN-based approaches (e.g., ChannelNet [9] and Channelformer [15]) in terms of channel estimation accuracy. In particular, the A-MMSE achieves approximately $43\%$ lower normalized MSE (NMSE) compared to conventional MMSE. We attribute these performance gains to the Attention Transformer’s ability to effectively capture nonstationary channel statistics, even under challenging high-mobility and frequency-selective fading conditions. Comparing to Channelformer, the A-MMSE achieves approximately $48\%$ lower NMSE. We note that these gain stem from our novel two-stage Attention encoder architecture, where frequency domain features are first extracted and subsequently refined through temporal correlation modeling. These channel estimation improvements directly translate to symbol detection performance gains, with A-MMSE achieving $13.1\%$ and $26.6\%$ lower bit error rate (BER) compared to MMSE and Channelformer respectively in the challenging mmWave scenario at $25$ dB SNR. Furthermore, the proposed method surpasses recent sensing-aided approaches [7], achieving up to a $35\%$ lower NMSE in high-mobility scenarios without the added complexity of echo signal processing. Moreover, the RA-A-MMSE further reduces computational complexity through adaptive rank reduction, maintaining over $99.6\%$ of the performance achieved by the full-rank A-MMSE. As a result, the RA-A-MMSE requires only $1.5\%$ of the computational complexity of Channelformer or MMSE (i.e., a $98.5\%$ reduction), while achieving approximately $36\%$ lower NMSE. The proposed methods also exhibit strong robustness to signal-to-ratio (SNR) mismatch, retaining nearly full performance ($\geq 98.1\%$) across a wide range of unseen SNR conditions. This underscores the practical efficiency and suitability of the proposed method for real-world 5G/6G implementations.
@@ III-A SP-based Approach
The SP-based OFDM channel estimation strategies have been extensively studied in [6].
@@
After obtaining the LS estimates at pilot positions, a subsequent interpolation (or extrapolation) step is typically applied to estimate the channel at non-pilot positions $(n,m)\in\mathcal{T}\setminus\mathcal{P}$. Various interpolation techniques can be employed, ranging from simple linear or bilinear interpolation [34, 4].
@@
where $\hat{\mathbf{H}}_{\text{init}}$ is the initial estimate (e.g., from LS or MMSE) and $\theta$ represents auxiliary information from environmental parameters [7] and channel statistics [8]. Enhancement functions include additive refinement ($\hat{\mathbf{H}}=\hat{\mathbf{H}}_{\text{init}}+\Delta\mathbf{H}(\theta)$), multiplication refinement ($\hat{\mathbf{H}}=\mathbf{W}(\theta)\hat{\mathbf{H}}_{\text{init}}$), and projection-based methods. While these approaches can improve estimation accuracy, they require additional sensing infrastructure and computational resources.
@@ III-B DNN-based Approach
ChannelNet: ChannelNet [9] represents one of the initial frameworks that exploits a DNN for OFDM channel estimation. In this approach, the channel response is treated as an image. An initial estimate of the channel is obtained by applying bilinear interpolation to sparsely located pilot symbols, resulting in a low-resolution and noise-contaminated channel map. Then, ChannelNet uses a super-resolution convolutional neural network (i.e., SRCNN) to recover high-resolution spatial features from the interpolated input. After this, a denoising convolutional neural network (i.e., DnCNN) is applied, incorporating residual learning and batch normalization to further refine the estimate.
Channelformer: Channelformer [15] exploits the Attention Transformer architecture for OFDM channel estimation. Unlike ChannelNet, Channelformer flattens the LS channel estimate for all pilot subcarriers into a single vector and feeds its real and imaginary parts as two separate input channels. Based on this, the encoder captures long-range dependencies among sparse pilot embeddings using the Attention Transformer architecture. Instead of the standard Feed-Forward Network, it employs convolutional layers to introduce local inductive bias, while retaining global modeling capabilities through multi-head self-Attention. The decoder refines these context-aware features using a residual convolutional network, integrating both global Attention and local structural information.
For better understanding, we explain the principle of Channelformer as follows. Fundamentally, Channelformer follows the Multi-Head Self-Attention (MHA) mechanism [13]. To be specific, denoting $p_{i}$ as the complex-valued channel observation at pilot index $i$ with $1\leq i\leq L$, each pilot observation is first embedded into a two-dimensional real-valued row vector that retains its inherent complex structure: $\mathbf{e}_{i}=\left[\Re(p_{i}),\Im(p_{i})\right]\in\mathbb{R}^{2}$. Then, we define the full sequence of embedded pilot vectors as
@@ IV-B1 Insights on Two-Stage Attention Encoder
Under the assumption of a wide-sense stationary uncorrelated scattering (WSSUS) environment and separable scattering function [35], the second-order statistics of OFDM channels exhibit a separable correlation function across frequency and temporal domains. Formally, the channel correlation matrix between two channel taps separated by frequency index difference $\Delta n$ and time index difference $\Delta m$ is factorized as follows:
@@
This two-stage encoder enables the model to learn the joint frequency–temporal correlation structure; each stage uses a single encoder block, which is critical for the accurate MMSE estimation. Crucially, we note that this encoder design represents a key distinction of the proposed A-MMSE learning mechanism compared to conventional DNN-based methods. Specifically, existing approaches typically adopt standard learning architectures such as CNN [9] or Attention Transformer [15], without explicitly tailoring them to the statistical properties inherent in OFDM channels. In contrast, our method explicitly incorporates the inherent temporal and frequency-domain correlation structures of OFDM channels into the learning process. This integration not only enhances learning efficiency but also improves the model’s capability to generalize under practical channel conditions. In this sense, the proposed A-MMSE approach is distinguished not only by its design, which enables the learning of a linear estimation filter rather than directly estimating the OFDM channel matrix (as illustrated in Fig.1), but also by its advanced learning mechanism that explicitly reflects the inherent correlation structure of OFDM channels.
@@ IV-B2 Multi-Head Self-Attention Mechanism
where $N_{\mathrm{RB}}$ is the number of resource blocks, $N_{\mathrm{SC/RB}}=12$ is the number of subcarriers per RB, and $N_{\mathrm{DM-RS\,sym}}$ denotes the number of OFDM symbols carrying DM-RS. The operation of the MHA module $\mathcal{F}_{\text{MHA}}(\cdot)$ is described in (13)–(16). We note that this allows each head to learn distinct frequency domain correlation patterns. Following the MHA module, Add & Layer Normalization is applied, succeeded by a FFN, and subsequently another Add & Layer Normalization, in sequential order. Details on Add & Norm and FFN are provided in [13]. After these processes, the encoder extracts frequency-domain statistical information, which can be formally represented as
@@ IV-B3 Training Method
where $N_{\mathcal{B}}$ denotes the number of mini batch and $\mathcal{B}$ denotes the mini batch size. In addition, $\mathbf{H}_{i,j}$ and $\hat{\bf{H}}_{i,j}$ represent the corresponding true channel matrix and estimated channel matrix, respectively. For the loss function $\mathcal{L}(\cdot)$, MSE or Huber loss can be chosen. We discuss this further in the next subsection. The loss function (32) is minimized via backpropagation and the Adam optimizer [36].

### v3

@@ I Introduction
Orthogonal frequency division multiplexing (OFDM) is a foundational waveform in modern wireless communication systems. OFDM was initially adopted in earlier standards such as Wi-Fi (IEEE 802.11), WiMAX, and was later introduced into the 3GPP-based cellular ecosystem with LTE. This adoption paved the way for OFDM’s continued prominence in 5G [shafi:jsac:17, guan:jsac:17] and its expected role in future 6G technologies [liy:twc:21].
@@
One of the critical problems in OFDM is channel estimation, which is essential for enabling coherent detection. The goal of OFDM channel estimation is to accurately estimate the channel state information for each subcarrier using noisy pilot symbols. A classical solution for the OFDM channel estimation involves signal processing (SP)-based approaches, including least square (LS) and minimum mean-squared error (MMSE) estimation methods [edfors:tcom:98, li:tvt:00]. LS is simple to implement, but its performance is limited because it cannot incorporate prior OFDM channel statistics. MMSE, in contrast, requires second-order statistics, such as noise variance and correlation across subcarriers, which are difficult to obtain accurately in practical scenarios [liu:survey:14].
@@
Beyond traditional SP-based approaches, sensing-aided channel estimation has emerged within the integrated sensing and communication (ISAC) frameworks [Qing2024IoTJ, Chen2023WCL]. These methods leverage communication signal echoes to extract sensing parameters, which then serve as prior information for channel refinement. Despite showing promise in exploiting communication-sensing synergy, these approaches require additional computational resources for sensing processing.
@@
Recently, inspired by the remarkable success of deep neural networks (DNNs) across various engineering fields, DNN-aided channel estimation for OFDM has gained considerable attention. DNNs, as pre-trained neural network models, can effectively capture complex data relationships without explicit prior statistical knowledge. Leveraging this data-driven approach, several prior works have explored DNN-aided OFDM channel estimation strategies. In [soltani:commlett:19], an OFDM channel slot was treated as image data. Based on this, a super-resolution network and a denoising neural network were simultaneously employed to estimate the OFDM channel (i.e., ChannelNet). In [li:wcl:20], using a similar modeling approach to [soltani:commlett:19], ReEsNet [li:wcl:20] was proposed, where a residual learning-based DNN was used to reconstruct the OFDM channel. This method achieves lower computational cost and higher accuracy than [soltani:commlett:19]. In [luan:wsa:21], a residual convolutional network combined with bilinear interpolation was proposed for OFDM channel estimation, offering enhanced flexibility to varying pilot patterns and reduced complexity. Beyond deterministic models, [balevi:jsac:21] optimized a deep generative model using compressive pilot measurements for high-dimensional channel estimation. This method leverages learned priors to reduce pilot overhead and surpass traditional sparse recovery techniques.
@@
In particular, with the advent of the Attention Transformer architecture [vaswani2017attention], some recent works investigated Transformer models as a compelling solution for wireless channel estimation problems, owing to their representational capacity. In [li2023wireless], super-resolution techniques were integrated with an Attention-based encoder-decoder framework to enhance OFDM channel estimation accuracy. In [luan2023channelformer], Channelformer was developed, where an Attention-based encoder-decoder neural architecture was designed to incorporate input precoding. A common structure in prior studies using DNNs for channel estimation [soltani:commlett:19, li:wcl:20, luan:wsa:21, balevi:jsac:21, li2023wireless, luan2023channelformer] is the use of a pre-trained neural network that directly maps noisy pilot signals to the estimated channel matrix, as illustrated in the left panel of Fig. 1.
@@
  • Computational overhead: Typically, DNN-based models involve a large number of parameters (approximately $53,000$ for ReEsNet [li:wcl:20] and $110,000$ for Channelformer [luan2023channelformer]), leading to significant computational and memory overhead per channel estimation. Furthermore, several nonlinear operations are used (e.g., ReLU [nair2010rectified], tanh [lecun2012efficient], ELU [clevert2015fast], and pooling [scherer2010evaluation]), limiting hardware acceleration effectiveness. These nonlinearities, although essential for model expressiveness, may introduce additional computational overhead and pose challenges related to inference complexity and practical implementation.
  • Limited interpretability: Existing DNN-based approaches are typically applied directly to OFDM channel estimation in a purely data-driven manner, without incorporating domain knowledge derived from mathematical models. However, studies such as [greenberg:neurips:23, ngu:wcommagg:23] showed that integrating domain-specific insights into neural network design can lead to notable performance gains over purely data-driven methods. This implies that considerable potential remains to further push performance boundaries by leveraging domain-specific insights.
@@
To address the challenges of purely data-driven approaches, recent research has actively explored the concept of model-based DNNs [mbdnn:book:23]. The core idea is to leverage a hybrid method that combines classical SP-based approaches and data-driven DNN-based approaches, thereby harnessing the advantages of both methodologies, i.e., powerful learning capability and mathematical interpretability. For instance, [kalmannet:tsp:22] proposed KalmanNet, where the Kalman gain matrix is learned using recurrent neural network architectures such as long short-term memory (LSTM) [lstm:97] or gated recurrent units (GRU) [cho:gru:14], while retaining the recursive update structure of the classical Kalman filtering algorithm [welch:kalman]. Thanks to its hybrid design, KalmanNet [kalmannet:tsp:22] effectively address the state estimation problem, even with partially known state models. Subsequently, several extensions were proposed, including [choi:tvt:23, shen2025kalmanformer, buch:tsp:24, aikalman:25]. In parallel, model-based DNN approaches have been widely applied to various tasks, such as multiple-input multiple-output (MIMO) beamforming design [xu:twc:25, pellaco:tsp:22], and MIMO detection [ngu:wcommagg:23, khani:twc:20]. However, to the best of our knowledge, no model-based DNN framework has been explored for OFDM channel estimation. This work addresses this notable gap by devising a novel channel estimation method that leverages not only the learning capability of DNNs but also domain knowledge from classical SP-based approaches.
@@
Via numerical simulations across diverse 3GPP TDL scenarios, we show that the proposed A-MMSE and RA-A-MMSE significantly outperform traditional SP-based approaches (e.g., LS and MMSE) and recent DNN-based approaches (e.g., ChannelNet [soltani:commlett:19] and Channelformer [luan2023channelformer]) in terms of channel estimation accuracy. In particular, the A-MMSE achieves approximately $43\%$ lower normalized MSE (NMSE) compared to conventional MMSE. We attribute these performance gains to the Attention Transformer’s ability to effectively capture nonstationary channel statistics, even under challenging high-mobility and frequency-selective fading conditions. Comparing to Channelformer, the A-MMSE achieves approximately $48\%$ lower NMSE. We note that these gain stem from our novel two-stage Attention encoder architecture, where frequency domain features are first extracted and subsequently refined through temporal correlation modeling. These channel estimation improvements directly translate to symbol detection performance gains, with A-MMSE achieving $13.1\%$ and $26.6\%$ lower bit error rate (BER) compared to MMSE and Channelformer respectively in the challenging mmWave scenario at $25$ dB SNR. Furthermore, the proposed method surpasses recent sensing-aided approaches [Qing2024IoTJ], achieving up to a $35\%$ lower NMSE in high-mobility scenarios without the added complexity of echo signal processing. Moreover, the RA-A-MMSE further reduces computational complexity through adaptive rank reduction, maintaining over $99.6\%$ of the performance achieved by the full-rank A-MMSE. As a result, the RA-A-MMSE requires only $1.5\%$ of the computational complexity of Channelformer or MMSE (i.e., a $98.5\%$ reduction), while achieving approximately $36\%$ lower NMSE. The proposed methods also exhibit strong robustness to signal-to-ratio (SNR) mismatch, retaining nearly full performance ($\geq 98.1\%$) across a wide range of unseen SNR conditions. This underscores the practical efficiency and suitability of the proposed method for real-world 5G/6G implementations.
@@ III-A SP-based Approach
The SP-based OFDM channel estimation strategies have been extensively studied in [liu:survey:14].
@@
After obtaining the LS estimates at pilot positions, a subsequent interpolation (or extrapolation) step is typically applied to estimate the channel at non-pilot positions $(n,m)\in\mathcal{T}\setminus\mathcal{P}$. Various interpolation techniques can be employed, ranging from simple linear or bilinear interpolation [coleri:tbc:02, edfors:tcom:98].
@@
where $\hat{\mathbf{H}}_{\text{init}}$ is the initial estimate (e.g., from LS or MMSE) and $\theta$ represents auxiliary information from environmental parameters [Qing2024IoTJ] and channel statistics [Chen2023WCL]. Enhancement functions include additive refinement ($\hat{\mathbf{H}}=\hat{\mathbf{H}}_{\text{init}}+\Delta\mathbf{H}(\theta)$), multiplication refinement ($\hat{\mathbf{H}}=\mathbf{W}(\theta)\hat{\mathbf{H}}_{\text{init}}$), and projection-based methods. While these approaches can improve estimation accuracy, they require additional sensing infrastructure and computational resources.
@@ III-B DNN-based Approach
ChannelNet: ChannelNet [soltani:commlett:19] represents one of the initial frameworks that exploits a DNN for OFDM channel estimation. In this approach, the channel response is treated as an image. An initial estimate of the channel is obtained by applying bilinear interpolation to sparsely located pilot symbols, resulting in a low-resolution and noise-contaminated channel map. Then, ChannelNet uses a super-resolution convolutional neural network (i.e., SRCNN) to recover high-resolution spatial features from the interpolated input. After this, a denoising convolutional neural network (i.e., DnCNN) is applied, incorporating residual learning and batch normalization to further refine the estimate.
Channelformer: Channelformer [luan2023channelformer] exploits the Attention Transformer architecture for OFDM channel estimation. Unlike ChannelNet, Channelformer flattens the LS channel estimate for all pilot subcarriers into a single vector and feeds its real and imaginary parts as two separate input channels. Based on this, the encoder captures long-range dependencies among sparse pilot embeddings using the Attention Transformer architecture. Instead of the standard Feed-Forward Network, it employs convolutional layers to introduce local inductive bias, while retaining global modeling capabilities through multi-head self-Attention. The decoder refines these context-aware features using a residual convolutional network, integrating both global Attention and local structural information.
For better understanding, we explain the principle of Channelformer as follows. Fundamentally, Channelformer follows the Multi-Head Self-Attention (MHA) mechanism [vaswani2017attention]. To be specific, denoting $p_{i}$ as the complex-valued channel observation at pilot index $i$ with $1\leq i\leq L$, each pilot observation is first embedded into a two-dimensional real-valued row vector that retains its inherent complex structure: $\mathbf{e}_{i}=\left[\Re(p_{i}),\Im(p_{i})\right]\in\mathbb{R}^{2}$. Then, we define the full sequence of embedded pilot vectors as
@@ IV-B1 Insights on Two-Stage Attention Encoder
Under the assumption of a wide-sense stationary uncorrelated scattering (WSSUS) environment and separable scattering function [Yoojin20152DLMMSE], the second-order statistics of OFDM channels exhibit a separable correlation function across frequency and temporal domains. Formally, the channel correlation matrix between two channel taps separated by frequency index difference $\Delta n$ and time index difference $\Delta m$ is factorized as follows:
@@
This two-stage encoder enables the model to learn the joint frequency–temporal correlation structure; each stage uses a single encoder block, which is critical for the accurate MMSE estimation. Crucially, we note that this encoder design represents a key distinction of the proposed A-MMSE learning mechanism compared to conventional DNN-based methods. Specifically, existing approaches typically adopt standard learning architectures such as CNN [soltani:commlett:19] or Attention Transformer [luan2023channelformer], without explicitly tailoring them to the statistical properties inherent in OFDM channels. In contrast, our method explicitly incorporates the inherent temporal and frequency-domain correlation structures of OFDM channels into the learning process. This integration not only enhances learning efficiency but also improves the model’s capability to generalize under practical channel conditions. In this sense, the proposed A-MMSE approach is distinguished not only by its design, which enables the learning of a linear estimation filter rather than directly estimating the OFDM channel matrix (as illustrated in Fig.1), but also by its advanced learning mechanism that explicitly reflects the inherent correlation structure of OFDM channels.
@@ IV-B2 Multi-Head Self-Attention Mechanism
where $N_{\mathrm{RB}}$ is the number of resource blocks, $N_{\mathrm{SC/RB}}=12$ is the number of subcarriers per RB, and $N_{\mathrm{DM-RS\,sym}}$ denotes the number of OFDM symbols carrying DM-RS. The operation of the MHA module $\mathcal{F}_{\text{MHA}}(\cdot)$ is described in (13)–(16). We note that this allows each head to learn distinct frequency domain correlation patterns. Following the MHA module, Add & Layer Normalization is applied, succeeded by a FFN, and subsequently another Add & Layer Normalization, in sequential order. Details on Add & Norm and FFN are provided in [vaswani2017attention]. After these processes, the encoder extracts frequency-domain statistical information, which can be formally represented as
@@ IV-B3 Training Method
where $N_{\mathcal{B}}$ denotes the number of mini batch and $\mathcal{B}$ denotes the mini batch size. In addition, $\mathbf{H}_{i,j}$ and $\hat{\bf{H}}_{i,j}$ represent the corresponding true channel matrix and estimated channel matrix, respectively. For the loss function $\mathcal{L}(\cdot)$, MSE or Huber loss can be chosen. We discuss this further in the next subsection. The loss function (32) is minimized via backpropagation and the Adam optimizer [kingma2014adam].